\begin{thebibliography}{10}

\bibitem{ALAGICRUSSELL}
Gorjan Alagic and Alexander Russell.
\newblock Uncertainty principles for compact groups.
\newblock {\em Illinois J. Math.}, 52(4):1315--1324, 2008.

\bibitem{ALIANTOINEGAZEAU}
S.~Twareque Ali, J.-P. Antoine, and J.-P. Gazeau.
\newblock Continuous frames in {H}ilbert space.
\newblock {\em Ann. Physics}, 222(1):1--37, 1993.

\bibitem{BANDEIRALEWISMIXON}
Afonso~S. Bandeira, Megan~E. Lewis, and Dustin~G. Mixon.
\newblock Discrete uncertainty principles and sparse signal processing.
\newblock {\em J. Fourier Anal. Appl.}, 24(4):935--956, 2018.

\bibitem{BORELLO}
Martino Borello, Wolfgang Willems, and Giovanni Zini.
\newblock On ideals in group algebras: an uncertainty principle and the {S}chur
  product.
\newblock {\em Forum Math.}, 34(5):1345--1354, 2022.

\bibitem{BOSSOLUCIANO}
Pasquale Bosso, Giuseppe~Gaetano Luciano, Luciano Petruzziello, and Fabian
  Wagner.
\newblock 30 years in: {\it quo vadis} generalized uncertainty principle?
\newblock {\em Classical Quantum Gravity}, 40(19):195014, 2023.

\bibitem{BUZANO}
Maria~Luisa Buzano.
\newblock Generalizzazione della diseguaglianza di {C}auchy-{S}chwarz.
\newblock {\em Rend. Sem. Mat. Univ. e Politec. Torino}, 31:405--409 (1974),
  1971/73.

\bibitem{CHUANGNG}
Kok~Seng Chua and Wee~Seng Ng.
\newblock A simple proof of the uncertainty principle for compact groups.
\newblock {\em Expo. Math.}, 23(2):147--150, 2005.

\bibitem{DEUTSCH}
David Deutsch.
\newblock Uncertainty in quantum measurements.
\newblock {\em Phys. Rev. Lett.}, 50(9):631--633, 1983.

\bibitem{DONOHOSTARK}
David~L. Donoho and Philip~B. Stark.
\newblock Uncertainty principles and signal recovery.
\newblock {\em SIAM J. Appl. Math.}, 49(3):906--931, 1989.

\bibitem{ELADBRUCKSTEIN}
Michael Elad and Alfred~M. Bruckstein.
\newblock A generalized uncertainty principle and sparse representation in
  pairs of bases.
\newblock {\em IEEE Trans. Inform. Theory}, 48(9):2558--2567, 2002.

\bibitem{EVRA}
Shai Evra, Emmanuel Kowalski, and Alexander Lubotzky.
\newblock Good cyclic codes and the uncertainty principle.
\newblock {\em Enseign. Math.}, 63(3-4):305--332, 2017.

\bibitem{FENG}
Tao Feng, Henk D.~L. Hollmann, and Qing Xiang.
\newblock The shift bound for abelian codes and generalizations of the
  {D}onoho-{S}tark uncertainty principle.
\newblock {\em IEEE Trans. Inform. Theory}, 65(8):4673--4682, 2019.

\bibitem{FOLLANDSITARAM}
Gerald~B. Folland and Alladi Sitaram.
\newblock The uncertainty principle: a mathematical survey.
\newblock {\em J. Fourier Anal. Appl.}, 3(3):207--238, 1997.

\bibitem{FFUJIIKUBO}
Masatoshi Fujii and Fumio Kubo.
\newblock Buzano's inequality and bounds for roots of algebraic equations.
\newblock {\em Proc. Amer. Math. Soc.}, 117(2):359--361, 1993.

\bibitem{GOHGOODMAN}
Say~Song Goh and Tim N.~T. Goodman.
\newblock Uncertainty principles in {B}anach spaces and signal recovery.
\newblock {\em J. Approx. Theory}, 143(1):26--35, 2006.

\bibitem{GOHMICCHELLI}
Say~Song Goh and Charles~A. Micchelli.
\newblock Uncertainty principles in {H}ilbert spaces.
\newblock {\em J. Fourier Anal. Appl.}, 8(4):335--373, 2002.

\bibitem{HEISENBERG}
W.~Heisenberg.
\newblock The physical content of quantum kinematics and mechanics.
\newblock In John~Archibald Wheeler and Wojciech~Hubert Zurek, editors, {\em
  Quantum Theory and Measurement}, Princeton Series in Physics, pages 62--84.
  Princeton University Press, Princeton, NJ, 1983.

\bibitem{JIANGLIUWU}
Chunlan Jiang, Zhengwei Liu, and Jinsong Wu.
\newblock Noncommutative uncertainty principles.
\newblock {\em J. Funct. Anal.}, 270(1):264--311, 2016.

\bibitem{KAISER}
Gerald Kaiser.
\newblock {\em A friendly guide to wavelets}.
\newblock Modern Birkh\"{a}user Classics. Birkh\"{a}user/Springer, New York,
  2011.

\bibitem{KEMPFMANGANOMANN}
Achim Kempf, Gianpiero Mangano, and Robert~B. Mann.
\newblock Hilbert space representation of the minimal length uncertainty
  relation.
\newblock {\em Phys. Rev. D (3)}, 52(2):1108--1118, 1995.

\bibitem{KRAUS}
K.~Kraus.
\newblock Complementary observables and uncertainty relations.
\newblock {\em Phys. Rev. D (3)}, 35(10):3070--3075, 1987.

\bibitem{KUPPINGER}
Patrick Kuppinger, Giuseppe Durisi, and Helmut B\"{o}lcskei.
\newblock Uncertainty relations and sparse signal recovery for pairs of general
  signal sets.
\newblock {\em IEEE Trans. Inform. Theory}, 58(1):263--277, 2012.

\bibitem{MAASSENUFFINK}
Hans Maassen and J.~B.~M. Uffink.
\newblock Generalized entropic uncertainty relations.
\newblock {\em Phys. Rev. Lett.}, 60(12):1103--1106, 1988.

\bibitem{MACCONEPATI}
Lorenzo Maccone and Arun~K. Pati.
\newblock Stronger uncertainty relations for all incompatible observables.
\newblock {\em Phys. Rev. Lett.}, 113(26):260401, 2014.

\bibitem{KRISHNA}
K.~Mahesh~Krishna.
\newblock Functional {D}eutsch uncertainty principle.
\newblock {\em J. Class. Anal.}, 23(1):11--18, 2024.

\bibitem{MESHULAM}
Roy Meshulam.
\newblock An uncertainty inequality for groups of order {$pq$}.
\newblock {\em European J. Combin.}, 13(5):401--407, 1992.

\bibitem{RICAUDTORRESANI}
Benjamin Ricaud and Bruno Torr\'{e}sani.
\newblock Refined support and entropic uncertainty inequalities.
\newblock {\em IEEE Trans. Inform. Theory}, 59(7):4272--4279, 2013.

\bibitem{ROBERTSON}
H.~P. Robertson.
\newblock The uncertainty principle.
\newblock {\em Phys. Rev.}, 34(1):163--164, 1929.

\bibitem{SCHRODINGER}
E.~Schr\"{o}dinger.
\newblock About {H}eisenberg uncertainty relation (original annotation by {A}.
  {A}ngelow and {M}.-{C}. {B}atoni).
\newblock {\em Bulgar. J. Phys.}, 26(5-6):193--203 (2000), 1999.
\newblock Translation of Proc. Prussian Acad. Sci. Phys. Math. Sect.
  {{\bf{1}}9} (1930), 296--303.

\bibitem{SELIG}
Kathi~K. Selig.
\newblock Uncertainty principles revisited.
\newblock {\em Electron. Trans. Numer. Anal.}, 14:165--177, 2002.

\bibitem{SMITH}
Kennan~T. Smith.
\newblock The uncertainty principle on groups.
\newblock {\em SIAM J. Appl. Math.}, 50(3):876--882, 1990.

\bibitem{STEELE}
J.~Michael Steele.
\newblock {\em The {C}auchy-{S}chwarz master class}.
\newblock Cambridge University Press, Cambridge, 2004.

\bibitem{STUDER}
Christoph Studer, Patrick Kuppinger, Graeme Pope, and Helmut B\"{o}lcskei.
\newblock Recovery of sparsely corrupted signals.
\newblock {\em IEEE Trans. Inform. Theory}, 58(5):3115--3130, 2012.

\bibitem{TAWFIKDIAB}
A.~Tawfik and A.~Diab.
\newblock Generalized uncertainty principle: Approaches and applications.
\newblock {\em International Journal of Modern Physics D}, 23(12):1430025,
  2014.

\bibitem{VONNEUMANNBOOK}
John von Neumann.
\newblock {\em Mathematical foundations of quantum mechanics}.
\newblock Princeton University Press, Princeton, NJ, 2018.

\bibitem{WIGDERSON}
Avi Wigderson and Yuval Wigderson.
\newblock The uncertainty principle: variations on a theme.
\newblock {\em Bull. Amer. Math. Soc. (N.S.)}, 58(2):225--261, 2021.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage{amssymb}
\usepackage{hyperref}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{definition}[theorem]{Definition}

\newtheorem{lemma} [theorem]{Lemma}

\newtheorem{problem}[theorem]{Problem}

\newtheorem{question}[theorem]{Question}
\title{This is the title}
\usepackage{fancyhdr}

\begin{document}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}	
\begin{center}
{\bf{Product Entropic Uncertainty Principle}}\\
\vspace{0.3cm}
\hrule\hrule\hrule\hrule\hrule
\vspace{0.3cm}
\textbf{K. Mahesh Krishna}\\
School of Mathematics and Natural Sciences\\
Chanakya University Global Campus\\
NH-648, Haraluru Village\\ 
Devanahalli Taluk, 	Bengaluru  Rural District\\
Karnataka  562 110, India\\
Email: kmaheshak@gmail.com\\

Date: \today
\end{center}

\hrule\hrule
\vspace{0.5cm}
\textbf{Abstract}: Motivated from Deutsch entropic uncertainty principle and several product uncertainty principles, we derive an uncertainty principle for the product of entropies using  functions.

\textbf{Keywords}:   Uncertainty Principle,  Frame.

\textbf{Mathematics Subject Classification (2020)}: 42C15.\\

\hrule

\hrule
\section{Introduction}
Let $\mathcal{H}$ be a finite dimensional Hilbert space. Given an orthonormal basis  $\{\omega_j\}_{j=1}^n$ for $\mathcal{H}$, the \textbf{Shannon entropy}  at a point $h \in \mathcal{H}_\tau$ is defined as 
\begin{align*}
	S_\tau (h)\coloneqq  \sum_{j=1}^{n} \left|\left \langle h, \tau_j\right\rangle \right|^2\log \left(\frac{1}{\left|\left \langle h, \tau_j\right\rangle \right|^2}\right)\geq 0, 
\end{align*}
where $\mathcal{H}_\tau \coloneqq \{h \in \mathcal{H}:\|h\|=1,  \langle h , \tau_j \rangle\neq 0, 1\leq j \leq n \}$. In 1983, Deutsch derived following uncertainty principle for Shannon entropy  \cite{DEUTSCH}.
\begin{theorem} (\textbf{Deutsch Uncertainty Principle}) \cite{DEUTSCH} \label{DU}
Let $\{\tau_j\}_{j=1}^n$,  $\{\omega_j\}_{j=1}^n$ be two orthonormal bases for a  finite dimensional Hilbert space $\mathcal{H}$. Then 
	\begin{align}\label{DUP}
2 \log n \geq S_\tau (h)+S_\omega (h)\geq -2 \log \left(\frac{1+\displaystyle \max_{1\leq j, k \leq n}|\langle\tau_j , \omega_k\rangle|}{2}\right)	\geq 0, \quad \forall h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega.
	\end{align}
\end{theorem}
Using Buzano inequality, it is easy to see that Theorem \ref{DU} holds for Parseval frames \cite{KRISHNA}. Note that  Inequality (\ref{DUP}) is for the sum of entropies and not for the product. It is best if we have uncertainty principles for the product because  once we have an uncertainty principle for the product, then the uncertainty principle for the sum follows from the AM-GM inequality. Following is a small list of uncertainty principle (UP) for the product.
\begin{enumerate}
	\item Heisenberg-Robertson-Schrodinger UP and its various generalizations \cite{SCHRODINGER, VONNEUMANNBOOK, ROBERTSON, HEISENBERG, FOLLANDSITARAM, SELIG, GOHMICCHELLI}.
	\item Donoho-Stark-Elad-Bruckstein-Ricaud-Torrésani support size UP \cite{ELADBRUCKSTEIN, RICAUDTORRESANI, DONOHOSTARK}.
	\item Smith UP and its generalizations \cite{SMITH, ALAGICRUSSELL, CHUANGNG, MESHULAM}.
	\item Kuppinger-Durisi-Bölcskei support size UP and its generalization \cite{KUPPINGER, STUDER}.
	\item A. Wigderson-Y. Wigderson UP \cite{WIGDERSON}.
	\item Maccone-Pati UP \cite{MACCONEPATI}.
	\item Goh-Goodman UP \cite{GOHGOODMAN}.
	\item Jiang-Liu-Wu subfactor UP \cite{JIANGLIUWU}.
\item Bandeira-Lewis-Mixon	numerical sparsity UP \cite{BANDEIRALEWISMIXON}.
	\item UP over finite fields \cite{EVRA, BORELLO, FENG}.
	\item Generalized UP \cite{BOSSOLUCIANO, TAWFIKDIAB, KEMPFMANGANOMANN}.
\end{enumerate}
	It seems that for logarithm functions, we can not have uncertainty for the product of entropies using coherence. In this paper,  we derive an uncertainty principle for  the product of entropies coming from certain functions.

\section{Product Entropic Uncertainty Principle}
In the paper,   $\mathbb{K}$ denotes $\mathbb{C}$ or $\mathbb{R}$ and $\mathcal{H}$  denotes a Hilbert space  (need not be finite dimensional) over $\mathbb{K}$.  We need the notion of continuous frame which  is  introduced independently by Ali, Antoine and Gazeau \cite{ALIANTOINEGAZEAU} and Kaiser \cite{KAISER}. 
\begin{definition}\cite{ALIANTOINEGAZEAU, KAISER}
	Let 	$(\Omega, \mu)$ be a measure space. A collection   $\{\tau_\alpha\}_{\alpha\in \Omega}$ in 	a  Hilbert  space $\mathcal{H}$ is said to be a \textbf{continuous Parseval frame}  for $\mathcal{H}$ if the following conditions hold.
	\begin{enumerate}[\upshape(i)]
		\item For each $h \in \mathcal{H}$, the map $\Omega \ni \alpha \mapsto \langle h, \tau_\alpha \rangle \in \mathbb{K}$ is measurable.
		\item 
		\begin{align*}
			\|h\|^2=\int\limits_{\Omega}|\langle h, \tau_\alpha \rangle|^2\,d\mu(\alpha), \quad \forall h \in \mathcal{H}.
		\end{align*}
	\end{enumerate}
\end{definition}
Recall that a continuous Parseval frame $\{\tau_\alpha\}_{\alpha\in \Omega}$   for $\mathcal{H}$ 	is said to be \textbf{1-bounded} if 
\begin{align*}
\|\tau_\alpha\|\leq 1, \quad \forall \alpha \in \Omega.
\end{align*}
Let $\phi:(0,1]\to (0, \infty)$ be a  function satisfying following.
\begin{enumerate}
	\item $\phi$ is continuous.
	\item $\phi$ is decreasing.
	\item $\phi(xy)\leq \phi(x)\phi(y)$, $ \forall x, y \in (0,1] $.
\end{enumerate}
In the entire paper, we assume that $\phi$ satisfies above conditions. Given such a $\phi$ and a 1-bounded continuous Parseval frame $\{\tau_\alpha\}_{\alpha\in \Omega}$   for $\mathcal{H}$, we define entropy 
	\begin{align*}
	S_{\tau,\phi}(h)=\int\limits_{\Omega}|\langle h, \tau_\alpha \rangle|^2\phi(|\langle h, \tau_\alpha \rangle|^2)\,d\mu(\alpha)\geq 0, \quad \forall h \in \mathcal{H}_\tau, 
\end{align*}
where $\mathcal{H}_\tau \coloneqq \{h \in \mathcal{H}:\|h\|=1,  \langle h , \tau_\alpha \rangle\neq 0, \forall \alpha \in \Omega \}$. Since $\phi$ is continuous, the integral exists. Before deriving the theorem   of this paper, we need a powerful inequality.
\begin{lemma} (\textbf{Buzano Inequality}) \cite{BUZANO, FFUJIIKUBO, STEELE} \label{BI}
	Let $\mathcal{H}$ be a Hilbert space. Then  
	\begin{align*}
		|\langle u, h \rangle \langle  h,v \rangle|\leq \|h\|^2\frac{\|u\|\|v\|+|\langle  u,v \rangle|}{2}, \quad \forall h, u, v \in \mathcal{H}.
	\end{align*}
\end{lemma}
\begin{theorem}\label{PD}
Let $(\Omega, \mu)$,  $(\Delta, \nu)$ be   measure spaces. Let $\{\tau_\alpha\}_{\alpha\in \Omega}$ and $\{\omega_\beta\}_{\beta\in \Delta}$ be 1-bounded 
	continuous Parseval frames   for $\mathcal{H}$. Then 
	\begin{align}\label{PDI}
	S_{\tau,\phi}(h)S_{\omega,\phi}(h)\geq 	\phi\left(\frac{\left(1+\sup_{\alpha\in \Omega, \beta\in \Delta}|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2}{4}\right)\geq 0, \quad \forall h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega.
	\end{align}
\end{theorem}
\begin{proof}
Let $h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega$. Then using the properties of $\phi$ and Lemma 	\ref{BI}, we get 
	\begin{align*}
		S_{\tau,\phi}(h)S_{\omega,\phi}(h)&=\left(\int\limits_{\Omega}|\langle h, \tau_\alpha \rangle|^2\phi(|\langle h, \tau_\alpha \rangle|^2)\,d\mu(\alpha)\right)\left(\int\limits_{\Delta}|\langle h, \omega_\beta \rangle|^2\phi(|\langle h, \omega_\beta \rangle|^2)\,d\nu(\beta)\right)\\
		&=\int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\phi(|\langle h, \tau_\alpha \rangle|^2)\phi(|\langle h, \omega_\beta \rangle|^2)\,d\nu(\beta)\,d\mu(\alpha)\\
		&\geq \int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\phi(|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2)\,d\nu(\beta)\,d\mu(\alpha)\\
		&\geq \int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\phi(|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2)\,d\nu(\beta)\,d\mu(\alpha)\\
		&\geq \int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\phi\left(\frac{\left(\|\tau_\alpha\|\omega_\beta\|+|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2}{4}\right)\,d\nu(\beta)\,d\mu(\alpha)\\
		&\geq \int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\phi\left(\frac{\left(1+\sup_{\alpha\in \Omega, \beta\in \Delta}|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2}{4}\right)\,d\nu(\beta)\,d\mu(\alpha)\\
		&=\phi\left(\frac{\left(1+\sup_{\alpha\in \Omega, \beta\in \Delta}|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2}{4}\right)\int\limits_{\Omega}\int\limits_{\Delta}|\langle h, \tau_\alpha \rangle|^2|\langle h, \omega_\beta \rangle|^2\,d\nu(\beta)\,d\mu(\alpha)\\
		&=\phi\left(\frac{\left(1+\sup_{\alpha\in \Omega, \beta\in \Delta}|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2}{4}\right).
	\end{align*}
\end{proof}
Theorem  \ref{PD}   gives the following question.
\begin{question}
	 Let $(\Omega, \mu)$,  $(\Delta, \nu)$ be   measure spaces,    $\mathcal{H}$ be a Hilbert space. For which pairs of 1-bounded continuous Parseval  frames $\{\tau_\alpha\}_{\alpha\in \Omega} $  and   $\{\omega_\beta\}_{\beta\in \Delta}$ for  $\mathcal{H}$, we have equality in Inequality (\ref{PDI})?
\end{question}
In 1988, Maassen  and Uffink (motivated from the conjecture  by Kraus made in 1987 \cite{KRAUS}) improved Deutsch entropic uncertainty principle.
\begin{theorem}\cite{MAASSENUFFINK} 
	(\textbf{Maassen-Uffink Entropic Uncertainty Principle})  \label{MU}
	Let $\{\tau_j\}_{j=1}^n$,  $\{\omega_k\}_{k=1}^n$ be two orthonormal bases for a  finite dimensional Hilbert space $\mathcal{H}$. Then 
	\begin{align*}
		S_\tau (h)+S_\omega (h)\geq -2 \log \left(\displaystyle \max_{1\leq j, k \leq n}|\langle\tau_j , \omega_k\rangle|\right), \quad \forall h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega.
	\end{align*}	
\end{theorem}
In 2013, Ricaud  and Torr\'{e}sani \cite{RICAUDTORRESANI} showed that orthonormal bases in Theorem \ref{MU} can be improved to Parseval frames. 
\begin{theorem}\cite{RICAUDTORRESANI} 
	(\textbf{Ricaud-Torr\'{e}sani Entropic Uncertainty Principle})  \label{RT}
	Let $\{\tau_j\}_{j=1}^n$,  $\{\omega_k\}_{k=1}^m$ be two Parseval frames  for a  finite dimensional Hilbert space $\mathcal{H}$. Then 
	\begin{align*}
		S_\tau (h)+S_\omega (h)\geq -2 \log \left(\displaystyle \max_{1\leq j \leq n, 1\leq k \leq m}|\langle\tau_j , \omega_k\rangle|\right), \quad \forall h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega.
	\end{align*}	
\end{theorem}
Proofs of  Theorems \ref{MU} and  \ref{RT} use Riesz-Thorin interpolation  and the differentiability   of logarithm function. We therefore  end by formulating the following problem. 
\begin{problem}
	Let $\{\tau_\alpha\}_{\alpha\in \Omega}$ and $\{\omega_\beta\}_{\beta\in \Delta}$ be 1-bounded 
	continuous Parseval frames   for $\mathcal{H}$. For which functions $\phi$, we have 
	\begin{align*}
		S_{\tau,\phi} (h)S_{\omega, \phi} (h)\geq   \phi\left(\left(\sup_{\alpha\in \Omega, \beta\in \Delta}|\langle \tau_\alpha, \omega_\beta \rangle|\right)^2\right), \quad \forall h \in \mathcal{H}_\tau \cap \mathcal{H}_\omega.
	\end{align*}		
\end{problem}

 \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}